# Brain Tumor Classification on MRI in Light of Molecular Markers

### v1

@@ Abstract
In research findings, co-deletion of the 1p/19q gene is associated with clinical outcomes in low-grade gliomas. The ability to predict 1p19q status is critical for treatment planning and patient follow-up. The aim of this study is to utilize a specially MRI-based convolutional neural network for brain cancers detection. Although public networks such as RestNet, AlexNet can effectively diagnose brain cancers using transfer learning, however, the model includes quite a few weights that have nothing to do with medical images. As a result, the diagnostic results are unreliable by transfer learning model. To deal with the problem of trustworthiness, we create the model from the ground up, rather than depending on a pre-trained model. To enable flexibility, we combined convolution stacking with a dropout and full connect operation, it improved performance by reducing overfitting. During model training, we also supplement the given dataset and inject Gaussian noise. We use three–fold cross-validation to train the best selection model. Comparing InceptionV3, VGG16, and MobileNetV2 fine-tuned with pre-trained models, our model produces better results. On an validation set of 125 codeletion vs. 31 not codeletion images, the proposed network achieves 96.37% percent F1-score, 97.46% percent precision, and 96.34% percent recall when classifying 1p/19q codeletion and not codeletion images.
@@ 1 Introduction
Tumors of the brain are malignant cell growths or aggregates in or around the brain.Depending on where they exist in the brain, low-grade gliomas[1, 2, 3] can manifest themselves in a variety of ways. The patient may have weakness or numbness in the right leg if the brain cancer grows in the area of the brain that controls it[1].
The most efficient method to detect brain tumors is MRI. Scanning generates a vast amount of magnetic resonance images, which is examined by a radiologist.Biomarker detection helps give patients the best appropriate treatment for their particular condition. This study is notable for the uniqueness and promising findings of merging deep learning with radiogenomics. Deep learning was better at detecting 1p/19q co-deletion on T2 images than it does on post-T1 contrast images.
@@ 4 Discussion
Small dataset can result in substantial training errors.In order to solve this problem, the training phase, Gaussian noise is included to improve the model’s normalizing ability and fault - tolerant. Adding noise means that the network cannot remember training samples because they are changing all the time, as a result of which network weights are reduced and the network becomes more robust with lower generalization error. Because fresh samples are drawn from the field proximity to existing samples, The input space’s shape has been smoothed out. It make easier for the network to learn the mapping function, resulting in better and faster learning.

### v2

@@ Abstract
In research findings, co-deletion of the 1p/19q gene is associated with clinical outcomes in low-grade gliomas. The ability to predict 1p19q status is critical for treatment planning and patient follow-up. This study aims to utilize a specially MRI-based convolutional neural network for brain cancer detection. Although public networks such as RestNet and AlexNet can effectively diagnose brain cancers using transfer learning, the model includes quite a few weights that have nothing to do with medical images. As a result, the diagnostic results are unreliable by the transfer learning model. To deal with the problem of trustworthiness, we create the model from the ground up, rather than depending on a pre-trained model. To enable flexibility, we combined convolution stacking with a dropout and full connect operation, it improved performance by reducing overfitting. During model training, we also supplement the given dataset and inject Gaussian noise. We use three–fold cross-validation to train the best selection model. Comparing InceptionV3, VGG16, and MobileNetV2 fine-tuned with pre-trained models, our model produces better results. On an validation set of 125 codeletion vs. 31 not codeletion images, the proposed network achieves 96.37% percent F1-score, 97.46% percent precision, and 96.34% percent recall when classifying 1p/19q codeletion and not codeletion images.
@@ 1 Introduction
Tumors of the brain are malignant cell growths or aggregates in or around the brain.Depending on where they exist in the brain, low-grade gliomas [2, 3] can manifest themselves in a variety of ways. The patient may have weakness or numbness in the right leg if the brain cancer grows in the area of the brain that controls it [1].
The most efficient method to detect brain tumors is MRI [17, 18]. Scanning generates a vast amount of magnetic resonance images, which is examined by a radiologist.Biomarker detection helps give patients the best appropriate treatment for their particular condition. This study is notable for the uniqueness and promising findings of merging deep learning with radiogenomics. Deep learning was better at detecting 1p/19q co-deletion on T2 images than it does on post-T1 contrast images.
@@ 4 Discussion
Small dataset [17] can result in substantial training errors.In order to solve this problem, the training phase, Gaussian noise is included to improve the model’s normalizing ability and fault - tolerant. Adding noise means that the network cannot remember training samples because they are changing all the time, as a result of which network weights are reduced and the network becomes more robust with lower generalization error. Because fresh samples are drawn from the field proximity to existing samples, The input space’s shape has been smoothed out. It make easier for the network to learn the mapping function, resulting in better and faster learning.